A father is 6 times as old as his son. Three years ago , the father was 9 times as old as his son. His find the present age of each
Answer by stanbon(75887)

A father is 6 times as old as his son. Three years ago , the father was 9 times as old as his son. His find the present age of each
-------------------------
Equations:
f = 6s
f-3 = 9(s-3)
----
Substitute for "f" and solve for "s"::
6s-3 = 9s - 27
3s = 24
s = 8 yrs (son's age)
--------
Solve for "f"::
f = 6s
f = 6*8 = 48 yrs (father's age)
---------------
Cheers,
Stan H.
------------